OpenStudy (anonymous):

If there are 4 Democrats and 3 Republicans on a committee, what is the probability that a Democrat will be the chair of the committee and a Republican will be the alternate? Assume that the chair and the alternate are chosen randomly. Round the answer to the nearest hundredth So confused it it 2 spots?

OpenStudy (anonymous):

and is it a combination?

OpenStudy (anonymous):

the probability that a democrat is chair is \(\frac{4}{7}\) then the probability that a republican is the alternate, given that the chair is a democrat, is \(\frac{3}{6}\) (since one person has already been picked, and it was a demcrat) then multiply to get your answer \[\frac{4}{7}\times\frac{1}{2}=\frac{2}{7}\]

OpenStudy (anonymous):

if you think of it as picking the alternate first, you get the same answer via \[\frac{3}{7}\times \frac{4}{6}\]

OpenStudy (anonymous):

Got it so I get that that you took 7 to 6 cause a democrat already had it thanks

OpenStudy (anonymous):

exactly you can also compute as \[\frac{4\times 3}{\binom{7}{2}}\] but that is more complicated to compute

OpenStudy (anonymous):

What about this A committee of 3 people is to be randomly selected from a group of 5 women and 8 men. What is the probability that the committee will consist only of women?

OpenStudy (anonymous):

5/8(3/7)?